Assume that you suspect that a dice has been tampered with. You throw the dice 36 times and record the number of times each value on the dice occurs in the table below. Conduct a hypothesis test to see if this dice is considered to be a fair dice (good dice and, therefore, has not been tampered with) at an α=0.01

Table 1.

Observed frequency (O) of the number of dots of die

Number of dots

Observed frequency

1

9

2

5

3

7

4

4

5

3

6

8

2.   A large big-box retailer is accused of sex discrimination in promotion decisions. The personnel files show there were 3300 male employees among them 960 were promoted in the past 6 months. During the same time, there were 2700 female employees, and among them, 700 were promoted. Do the data substantiate the claim that this large big-box retailer discriminated against female workers in promotion decisions? Conduct a hypothesis test to evaluate if there is a relationship between promotion decisions and employee’s sex at an α=0.05.

Table 2.

Observed frequency (O) between promotion decision and employee’s sex

Promoted

Not promoted

Row total

Male

960

2340

3300

Female

700

2000

2700

Column total

1660

4340

6000
Answer #1

(1)

H0: Null Hypothesis: This dice is considered to be a fair dice (good dice and, therefore, has not been tampered with) (Claim)

HA: Alternative Hypothesis: This dice is considered to be not a fair dice (bad dice and, therefore, has been tampered with)

Test Statistic (\chi ^{2}) is got asfollows:

Observed (O) Expected (E) (O - E)2/E
9 6 1.500
5 6 0.167
7 6 0.167
4 6 0.667
3 6 1.500
8 6 0.667
Total = \chi ^{2} = 4.668

Degrees of Fredom =6 - 1 = 5

From Table, critical value of \chi ^{2} = 15.086

Since calculated value of \chi ^{2} = 4.668 is less than critical value of \chi ^{2} = 15.086, the difference is not significant. Fail to reject the null hypothesis.

Conclusion:
The data support the claim that this dice is considered to be a fair dice (good dice and, therefore, has not been tampered with).

(2)
H0: Null Hypothesis: There is no relationship between promotion decisions and employee’s sex

HA: Alternative Hypothesis: There is a relationship between promotion decisions and employee’s sex (Claim)

Expected Frequencies are got as follows:

Promoted Not promoted Row total
Male 1660 X3300/6000 = 913.00 4340 X3300/6000 = 2387.00 3300
Female 1660 X2700/6000 = 747.00 4340 X2700/6000 = 1953.00 2700

Column total

1660

4340 6000

Test Statistc (\chi ^{2}) is got as follows:

Observed (O) Expected (E) (O - E)2/E
960 913 2.42
230 2387 0.93
700 747 2.96
2000 1953 1.13
Total = \chi ^{2} = 7.43

df = (r - 1) X(c - 1) = (2 - 1) X(2 - 1) = 1

By Technology, p - value = 0.0064

Since p - value = 0.0064 is less than α=0.05, the differenc is significant. Reject null hypothesis.

Conclusion:
The data support the claim that there is a relationship between promotion decisions and employee’s sex